# Effort: Efficient Orthogonal Modeling for Generalizable AI-Generated Image Detection

### v1

@@ C.1 Additional Ablation Studies
Appendix D Additional Analysis and Visualizations
D.1 Logit Distribution Analysis
We show the prediction distribution of a naive trained detector’s output (real probabilities) in Fig. 9. We see that the vanilla CNN detector tends to learn the fake in the training set and cannot learn a good feature of real when evaluating unseen fakes. When the fake patterns are unseen during training, the logit for real/fake classification tends to be very unconfident. Additionally, the overlapping between real and fake classes of Xception is significantly increased when detecting unseen the forgery dataset. In contrast, our approach preserves semantic knowledge while learning forgery patterns during training, thereby improving the model’s generalization performance.
Figure 10: t-SNE visualizations of the latent feature spaces between vanilla CNN [86] and ours. We use the testing set of ProGAN and StyleGAN within UniversalFakeDetect Dataset [86] for visualization.
Figure 11: t-SNE visualizations of the latent feature spaces between vanilla CNN [86] and ours. We use the testing set of CycleGAN within UniversalFakeDetect Dataset [86] for visualization.
D.2 Additional t-SNE Visualizations
In our manuscript, we visualize the t-SNE of the seen fake ProGAN and unseen fake StyleGAN for the comparison of vanilla CNN (Res-50 [86]) and ours (see Fig. 10). Now, we provide additional t-SNE visualizations, including t-SNE results between vanilla CNN and ours on CycleGAN, as shown in Fig. 11. As we can see from both two figures, our approach maximizes and preserves the semantic knowledge while fitting the forgery patterns during training, whereas the vanilla CNN overfits the seen fake method, learning forgery patterns only, thereby resulting in a highly low-ranked feature space (see Fig. 6 and Fig. 7 of the manuscript for details) and causing the overfitting to seen forgery patterns in the training set. Additionally, we see that the logit distribution of the vanilla CNN has a larger overlapping between fake and real, while ours is highly smaller, suggesting that our approach achieves a better generalization performance.
Appendix E Limitation and Future Work
The core idea of this paper is to decompose the original semantic space into two orthogonal subspaces for preserving semantic knowledge while learning the forgery. In our manuscript and appendix, we have conducted extensive experiments and in-depth analysis on both deepfake and synthetic image detection benchmarks, showing the superior advantages in both generalization and efficiency. One limitation of our work is that our approach regards all fake methods in one class during training real/fake classifiers, potentially ignoring the specificity and generality of different fake methods.

### v2

@@ D.1 Additional Ablation Studies
Appendix E Additional Analysis and Visualizations
E.1 Additional t-SNE Visualizations
In our manuscript, we visualize the t-SNE of the seen fake ProGAN and unseen fake StyleGAN for the comparison of vanilla CNN (Res-50 (Wang et al., 2020b)) and ours (see Fig. 9). Now, we provide additional t-SNE visualizations, including t-SNE results between vanilla CNN and ours on CycleGAN, as shown in Fig. 10. As we can see from both two figures, our approach maximizes and preserves the pre-trained knowledge while fitting the forgery patterns during training, whereas the vanilla CNN overfits the seen fake method, learning forgery patterns only, thereby resulting in a highly low-ranked feature space (see Fig. 6 and Fig. 7 of the manuscript for details) and causing the overfitting to seen forgery patterns in the training set. Additionally, we see that the logit distribution of the vanilla CNN has a larger overlapping between fake and real, while ours is highly smaller, suggesting that our approach achieves a better generalization performance.
Appendix F Limitation and Future Work
The core idea of this paper is to decompose the original feature space into two orthogonal subspaces for preserving pre-trained knowledge while learning the forgery. In our manuscript and supplementary, we have conducted extensive experiments and in-depth analysis on both deepfake and synthetic image detection benchmarks, showing the superior advantages in both generalization and efficiency. One limitation of our work is that our approach regards all fake methods in one class during training real/fake classifiers, potentially ignoring the specificity and generality of different fake methods.